# Is It Navajo? Accurate Language Detection in Endangered Athabaskan Languages

### v1

@@ Abstract
Endangered languages, such as Navajo—the most widely spoken Native American language—are significantly underrepresented in contemporary language technologies, exacerbating the challenges of their preservation and revitalization. This study evaluates Google’s large language model (LLM)-based language identification system, which consistently misidentifies Navajo, exposing inherent limitations when applied to low-resource Native American languages. To address this, we introduce a random forest classifier trained on Navajo and eight frequently confused languages. Despite its simplicity, the classifier achieves near-perfect accuracy (97-100%), significantly outperforming Google’s LLM-based system. Additionally, the model demonstrates robustness across other Athabaskan languages—a family of Native American languages spoken primarily in Alaska, the Pacific Northwest, and parts of the Southwestern United States—suggesting its potential for broader application. Our findings underscore the pressing need for NLP systems that prioritize linguistic diversity and adaptability over centralized, one-size-fits-all solutions, especially in supporting underrepresented languages in a multicultural world. This work directly contributes to ongoing efforts to address cultural biases in language models and advocates for the development of culturally localized NLP tools that serve diverse linguistic communities.
@@
Ivory Yang1  Weicheng Ma2  Chunhui Zhang3  Soroush Vosoughi4 1,2,3,4Department of Computer Science, Dartmouth College {Ivory.Yang.GR, Weicheng.Ma, Chunhui.Zhang.gr, Soroush.Vosoughi}@dartmouth.edu
@@ 1 Introduction
The urgency of preserving endangered languages is not merely a linguistic issue but one deeply connected to the preservation of cultural, historical, and ecological knowledge Tulloch (2006); Zariquiey et al. (2022); Zhang et al. (2022); Cusenza and Çöltekin (2024). These languages reflect the intellectual heritage of diverse communities, playing a critical role in maintaining global cultural diversity. Yet, despite this significance, the development of language technologies has been disproportionately focused on languages with large speaker bases and economic clout, leaving languages with smaller populations—such as Native American languages—largely unsupported.
This study centers on Navajo, the most widely spoken Native American language Dietrich et al. (2022), which remains vastly underrepresented in widely-used digital platforms like Google Translate, powered by large language models (LLMs) (Google Cloud, 2024). The lack of comprehensive linguistic datasets and dedicated tools impedes both language preservation and learning efforts Shamsfard (2019). This technological gap is even more pronounced for other Native American languages, many of which are on the verge of extinction due to minimal technological integration and educational resources Meredith (2013); Flavelle and Lachler (2023).
Figure 1: Recognizing Navajo (and other Athabaskan languages) presents significant challenges for centralized models like Google Translate. Our model addresses these challenges effectively.
Our investigation into Google Translate’s performance on the Navajo 10k dataset Goldhahn et al. (2012) revealed consistent misidentification of Navajo as unrelated languages, exposing significant shortcomings in its algorithmic treatment of Native American languages. In response, we developed a language identification model tailored to accurately distinguish Navajo from frequently misidentified languages, achieving near-perfect accuracy. This success illustrates that low-resource languages, often overlooked by major technological platforms, can be effectively supported with targeted approaches and resources. Beyond Navajo, we extended our model to other languages in the Athabaskan family—including Western Apache, Mescalero Apache, Jicarilla Apache, and Lipan Apache George and Lopraisová . Our model’s robustness across these related languages underscores its potential applicability across broader linguistic groups (see Figure 1). This suggests a viable path for NLP technologies to not only support individual endangered languages but to facilitate revitalization efforts across entire language families.
@@ 2 Background
While there are studies on endangered languages Zariquiey et al. (2022); Zhang et al. (2022); Cusenza and Çöltekin (2024), their integration into business technologies remains insufficient. For example, although Google Translate supports over a hundred languages, it fails to include Navajo, often providing inaccurate or irrelevant suggestions. This gap highlights the limited support for Native American languages. Similarly, advanced NLP models, such as ChatGPT, struggle with Navajo due to a lack of training data, which is predominantly derived from more widely spoken languages Hangya et al. (2022).
@@ 3 Identification for NatAm Languages
Figure 2: Misdetected languages by Google Language Detection API, along with their frequency counts.
The benefits of Native American language identification (especially for underrepresented ones such as Navajo) are twofold: to bolster the development of linguistic tools tailored to these languages, and to aid in their preservation and revitalization Mohanty et al. (2023). Effective identification is foundational for creating technologies that understand and process these languages, addressing the significant digital divide in language technology support Mohanty et al. (2024).
Our evaluation aims at a detailed assessment of the models’ capability to accurately recognize and differentiate between Native American languages and others. By understanding the strengths and limitations of our models, we can refine our techniques to better serve the needs of Native American language communities.
@@ 3.1 Benchmark Construction
To construct our dataset for evaluating language identification models, we used two distinct approaches to account for the diversity and specificity needed to test our models thoroughly. The first dataset was formed based on eight languages that Google Translate frequently misidentifies as Navajo. This subset includes samples from several unrelated languages that Google’s algorithm erroneously suggests as similar, namely Icelandic, Lingala, Wolof, Czech, Polish, Manx, Fulah and Yoruba, with their distribution shown in Figure 2. Each entry consists of 1k or 10k sentences extracted from the Leipzig corpora Goldhahn et al. (2012), ensuring standardization against the Navajo 10k dataset in terms of linguistic features and contexts.
@@ 3.3 Experimental Results
Figure 3: Our solution’s classification results for nine languages. The x-axis is the predicted language classes, while the y-axis is the true language classes. The diagonal elements represent the number of correctly classified instances for each language, while off-diagonal elements indicate misclassifications. The intensity of the color indicates the frequency of classifications. Labels on x and y axes correspond to languages in Table 1.
Language Class Precision Recall F1-score Support
Navajo 0 0.97 1.00 0.98 1980
Icelandic 1 0.99 0.99 0.99 1989
Lingala 2 0.99 0.97 0.98 265
Wolof 3 1.00 0.99 0.99 2023
Czech 4 0.97 0.95 0.96 2028
Polish 5 0.96 0.98 0.97 1970
Manx 6 0.99 0.99 0.99 1974
Fulah 7 0.96 0.92 0.94 219
Yoruba 8 0.99 0.98 0.99 2010
Our experimental evaluation of the Random Forest classifier is conducted using a dataset that included Navajo, labeled as ’0’, alongside a selection of languages that Google Translate frequently misidentifies as similar to Navajo. The dataset comprised 57,832 training samples and 14,458 test samples, each vectorized into 5,000 features to capture a wide range of linguistic attributes. The classifier achieved an overall accuracy of 98%, demonstrating its strong capability to distinguish Navajo from the misidentified languages.
The performance metrics show high precision, recall, and f1-scores across all tested languages, with Navajo achieving a precision of 0.97, a recall of 1.00, and an f1-score of 0.98, as shown in Table 1. These results highlight the classifier’s effectiveness in accurately differentiating Navajo from other unrelated languages, which are often erroneously suggested by Google Translate. The confusion matrix in Figure 3 highlights the classifier’s performance, particularly for Navajo. The model achieved 1,976 true positives for Navajo, with only a few false negatives. These results demonstrate the model’s robustness in identifying Navajo and underscore its potential for supporting language identification in underrepresented Native American languages. Nevertheless, addressing the few misclassifications through enhanced training could further improve accuracy and generalization.
The stable performance of our classifier reveals a flaw in Google Translate’s language identification: its lack of support for Native American languages, including Navajo. This absence contributes to the erroneous suggestions of linguistically unrelated languages, highlighting a critical need to include these languages in global technology platforms to better respect and reflect linguistic diversity.
@@ 3.4 Model Generalizability
Following the successful differentiation of Navajo from languages inaccurately suggested by Google Translate, we further tested the classifier’s capability with our second dataset of curated Apache languages. Upon running this subset through the classifier, initially trained to distinguish Navajo from other languages, we observed that the classifier often identified these Apache languages as Navajo. This result is particularly significant given the linguistic similarities within the Athabaskan language family, to which both Navajo and the Apache languages belong. The classifier’s performance here underscores its ability not only to identify Navajo with high accuracy but also to generalize across related languages within the same family. This generalizability is indicative of the model’s potential utility in broader linguistic applications, especially in creating tools that support multiple but related Native American languages.
These findings also raise interesting questions about the classifier’s sensitivity to the nuances between closely related languages and its potential role in developing more sophisticated NLP tools that can accurately differentiate between languages with subtle linguistic differences. The detection for Navajo performed best for Western Apache and Mescalero Apache, as shown in Table 2. Both these languages fall under the Western Apachean subgroup along with Navajo, as shown in Figure 4. On the other hand, Jicarilla Apache and in particular, Lipan Apache, performed less well in Navajo detection, which could be because they fall under the Eastern Apachean subgroup. This could be pivotal for linguistic preservation, allowing for the development of specialized educational and communicational tools tailored to each language’s unique characteristics.
@@ 4 Conclusion and Future Work
This study demonstrates the effectiveness of our Random Forest classifier in accurately distinguishing Navajo from languages commonly misidentified by Google Translate, as well as effectively recognising related Athabaskan languages. These results highlight the potential for broader applications in language identification, particularly for underrepresented languages. Our findings highlight a significant gap in support for Native American languages in current digital platforms, and highlight the need for refined, inclusive language models. This work illustrates the promise of using machine learning to improve digital accessibility and usability for endangered languages, thereby contributing to language preservation efforts.
Future work can focus on expanding the classifier’s training to include additional Native American languages, improving its adaptability, and extending its utility to different language groups. The development of tools capable of distinguishing closely related languages is crucial for supporting educational and communication needs within Native communities. We also advocate the decentralisation of NLP research efforts, emphasising the need for targeted investment in endangered languages. Such initiatives are essential to ensure that advances in language technology promote linguistic equity, thereby preserving cultural diversity and heritage in the digital age.
@@
While the study successfully demonstrates the Random Forest classifier’s efficacy in distinguishing Navajo from languages commonly misidentified by Google Translate and identifying related Athabaskan languages, it does have limitations that impact its broader applicability. Firstly, the language variety included in the study is limited; the classifier was tested primarily against a small set of languages suggested by Google and a few Athabaskan languages. This narrow scope might not capture the classifier’s effectiveness across a broader range of Native American languages, potentially limiting its utility for other endangered language families. Secondly, the experimental design assumes a binary distinction between Navajo and other languages without considering the intra-group variations and dialectical differences that exist within the Athabaskan language family, which could affect the classifier’s accuracy in real-world applications. Lastly, the reliance on vectorized features of 5,000 dimensions may overlook some finer linguistic nuances, which are crucial for distinguishing between closely related languages. Addressing these limitations in future work will be essential for developing more robust and widely applicable language identification systems.
@@
Ethical considerations are paramount in the development of language technology, especially for Native American languages, which are deeply intertwined with cultural identity and heritage. This study emphasizes the importance of respectful engagement with these communities, recognizing the cultural, spiritual, and historical significance of their languages. Technology development involving Native American languages should proceed with close collaboration with native speakers and community leaders to ensure that these tools support and reinforce language preservation rather than contributing to cultural homogenization or appropriation. Additionally, data privacy and consent are critical, as much of the linguistic data involves sensitive cultural content. Ensuring that communities retain control over how their linguistic resources are used is essential for maintaining trust and upholding ethical standards in research. Moreover, to ensure transparency and foster research, we have made our code and datasets used publicly available through an anonymous GitHub (https://github.com/isitnavajo/NavajoDetector).
@@ Appendix A Aligned Words Across Native American Languages
Figure 5: Aligned Words from Chiricahua Apache Mimbreno Nde Nation (2024).

### v2

@@ Abstract
Endangered languages, such as Navajo—the most widely spoken Native American language—are significantly underrepresented in contemporary language technologies, exacerbating the challenges of their preservation and revitalization. This study evaluates Google’s Language Identification (LangID) tool, which does not currently support any Native American languages. To address this, we introduce a random forest classifier trained on Navajo and twenty erroneously suggested languages by LangID. Despite its simplicity, the classifier achieves near-perfect accuracy (97-100%). Additionally, the model demonstrates robustness across other Athabaskan languages—a family of Native American languages spoken primarily in Alaska, the Pacific Northwest, and parts of the Southwestern United States—suggesting its potential for broader application. Our findings underscore the pressing need for NLP systems that prioritize linguistic diversity and adaptability over centralized, one-size-fits-all solutions, especially in supporting underrepresented languages in a multicultural world. This work directly contributes to ongoing efforts to address cultural biases in language models and advocates for the development of culturally localized NLP tools that serve diverse linguistic communities.
@@
Ivory Yang  Weicheng Ma  Chunhui Zhang  Soroush Vosoughi Department of Computer Science, Dartmouth College {Ivory.Yang.GR, Weicheng.Ma, Chunhui.Zhang.GR, Soroush.Vosoughi}@dartmouth.edu
@@ 1 Introduction
The urgency of preserving endangered languages is not merely a linguistic issue but one deeply connected to the preservation of cultural, historical, and ecological knowledge Tulloch (2006); Zariquiey et al. (2022); Zhang et al. (2022); Cusenza and Çöltekin (2024); Yang et al. (2025). These languages reflect the intellectual heritage of diverse communities, playing a critical role in maintaining global cultural diversity. Yet, despite this significance, the development of language technologies has been disproportionately focused on languages with large speaker bases and economic clout, leaving languages with smaller populations—such as Native American languages—largely unsupported.
This study centers on Navajo, the most widely spoken Native American language Dietrich et al. (2022), which remains unsupported by commercial language technologies like Google’s Language Identification (LangID) tool Caswell et al. (2020). The lack of comprehensive linguistic datasets and dedicated tools impedes both language preservation and learning efforts Shamsfard (2019). This technological gap is even more pronounced for other Native American languages, many of which are on the verge of extinction due to minimal technological integration and educational resources Meredith (2013); Flavelle and Lachler (2023).
Figure 1: Google’s LangID does not currently support any Native American languages, including Navajo and other Athabaskan languages. Our model addresses these challenges effectively.
Google LangID’s performance on the Navajo 10k dataset Goldhahn et al. (2012) revealed complete misidentification of Navajo sentences as unrelated languages, an expected outcome given that LangID does not currently support any Native American language. In response, we developed a language identification model tailored to accurately distinguish Navajo from languages erroneously suggested by LangID, achieving near-perfect accuracy. This success illustrates that low-resource languages, often overlooked by major technological platforms, can be effectively supported with targeted approaches and resources. Beyond Navajo, we extended our model to other languages in the Athabaskan family—including Western Apache, Mescalero Apache, Jicarilla Apache, and Lipan Apache George and Lopraisová . Our model’s robustness across these related languages underscores its potential applicability across broader linguistic groups (see Figure 1). This suggests a viable path for NLP technologies to not only support individual endangered languages but to facilitate revitalization efforts across entire language families.
@@ 2 Background
While there exist studies on endangered languages Zariquiey et al. (2022); Zhang et al. (2022); Cusenza and Çöltekin (2024), their integration into business technologies remains insufficient. For example, although Google’s LangID supports over a hundred languages, it fails to include any Native American language, and so provides completely inaccurate suggestions when encountering Navajo. Similarly, advanced NLP models, such as ChatGPT, struggle with Navajo due to a lack of training data, which is predominantly derived from more widely spoken languages Hangya et al. (2022).
@@ 3 Identification for NatAm Languages
Figure 2: Visual representation of some erroneously suggested languages by Google LangID, along with their frequency counts.
The benefits of Native American language identification are twofold: to bolster the development of linguistic tools tailored to these languages, and to aid in their preservation and revitalization Mohanty et al. (2023). Effective identification is foundational for creating technologies that understand and process these languages, addressing the significant digital divide in language technology support Mohanty et al. (2024). Our evaluation aims at a detailed assessment of the models’ capability to accurately recognize and differentiate between Native American languages and others. By understanding the strengths and limitations of our models, we can refine our techniques to better serve the needs of Native American language communities.
@@ 3.1 Benchmark Construction
To construct our dataset for evaluating language identification models, we used two distinct approaches to account for diversity and specificity. The first dataset was formed based on twenty languages [^1] The languages are Icelandic, Lingala, Wolof, Czech, Polish, Manx, Fulah, Yoruba, Portuguese, Somali, Slovak, Tsonga, Spanish, Oromo, Indonesian, Igbo, Northern Sami, Irish, Arabic and English. that Google’s LangID misidentifies as when presented with Navajo sentences, with their distribution shown in Figure 2. Each entry consists of 1k or 10k sentences extracted from the Leipzig corpora Goldhahn et al. (2012), ensuring standardization against the Navajo 10k dataset in terms of linguistic features and contexts.
@@
Figure 3: Classification results for Navajo and 20 other languages, presented as a confusion matrix. The intensity of the color indicates the frequency of classifications. Labels on x and y axes correspond to the languages listed in Table 1.
@@ 3.2 Model Specifics
Language Class Precision Recall F1-score Support
Navajo 0 1.00 0.99 0.99 1975
Icelandic 1 0.99 0.99 0.99 1971
Lingala 2 0.94 0.95 0.94 257
Wolof 3 1.00 0.98 0.99 2014
Polish 4 0.91 0.95 0.93 2032
Czech 5 0.89 0.85 0.87 1987
Manx 6 0.99 0.99 0.99 1917
Fulah 7 0.92 0.88 0.90 202
Yoruba 8 0.99 0.96 0.97 2020
Portuguese 9 0.98 0.98 0.98 2030
Somali 10 0.98 0.99 0.99 1963
Slovak 11 0.88 0.88 0.88 2018
Tsonga 12 0.99 0.98 0.98 1929
Spanish 13 0.99 0.98 0.99 1979
Oromo 14 0.99 0.98 0.99 1980
Indonesian 15 0.99 0.98 0.99 2031
Igbo 16 0.99 0.96 0.98 2059
Northern Sami 17 0.98 0.97 0.98 1996
Irish 18 0.99 0.97 0.98 2046
Arabic 19 0.89 0.99 0.94 1998
English 20 0.96 0.98 0.97 2054
@@ 3.3 Experimental Results
Our experimental evaluation of the Random Forest classifier is conducted using a dataset that included Navajo, labeled as ’0’, alongside a selection of languages that Google’s LangID erroneously identifies as similar to Navajo. The dataset comprised 153,832 training samples and 38,458 test samples, each vectorized into 5,000 features to capture a wide range of linguistic attributes. The classifier achieved an overall accuracy of 97%, demonstrating its strong capability to distinguish Navajo from the misidentified languages.
The performance metrics show high precision, recall, and f1-scores across all tested languages, with Navajo achieving a precision of 1.00, a recall of 0.99, and an f1-score of 0.99, as shown in Table 1. These results highlight the classifier’s effectiveness in accurately differentiating Navajo from other unrelated languages, which are erroneously suggested by Google’s LangID. The confusion matrix in Figure 3 highlights the classifier’s performance, particularly for Navajo. The model achieved 1,946 true positives for Navajo, with only a few false negatives. These results demonstrate the model’s robustness in identifying Navajo and underscore its potential for supporting language identification in underrepresented Native American languages. Nevertheless, addressing the few misclassifications through enhanced training could further improve accuracy and generalization.
The stable performance of our classifier serves as a counterresponse to Google’s LangID, and its lack of support for Native American languages, including Navajo. This omission directly results in the erroneous suggestions of linguistically unrelated languages, highlighting a critical need to include these languages in global technology platforms to better respect and reflect linguistic diversity.
@@ 3.4 Model Generalizability
Following the successful differentiation of Navajo from languages erroneously suggested by Google’s LangID, we further tested the classifier’s capability with our second dataset of curated Apache languages. Upon running this subset through the classifier, initially trained to distinguish Navajo from other languages, we observed that the classifier often identified these Apache languages as Navajo. This result is particularly significant given the linguistic similarities within the Athabaskan language family, to which both Navajo and the Apache languages belong. The classifier’s performance here underscores its ability not only to identify Navajo with high accuracy but also to generalize across related languages within the same family. This generalizability is indicative of the model’s potential utility in broader linguistic applications, especially in creating tools that support multiple but related Native American languages.
These findings also raise interesting questions about the classifier’s sensitivity to the nuances between closely related languages and its potential role in developing more sophisticated NLP tools that can accurately differentiate between languages with subtle linguistic differences. The detection for Navajo performed best for Western Apache and Mescalero Apache, as shown in Table 2. Both these languages fall under the Western Apachean subgroup along with Navajo, as shown in Figure 4. On the other hand, Jicarilla Apache and in particular, Lipan Apache, performed less well in Navajo detection, which could be because they fall under the Eastern Apachean subgroup. This observation could be pivotal for linguistic preservation, allowing for the development of specialized educational and communicational tools tailored to each language’s unique characteristics.
@@ 4 Conclusion and Future Work
This study demonstrates the effectiveness of our Random Forest classifier in accurately distinguishing Navajo from languages erroneously suggested by Google’s LangID, as well as effectively recognizing related Athabaskan languages. These results emphasize the potential for broader applications in language identification, particularly for underrepresented languages. Our findings highlight a significant gap in support for Native American languages in current digital platforms, and urge the need for refined, inclusive language models.
Future work can focus on expanding the classifier’s training to include additional Native American languages, improving its adaptability, and extending its utility to different language groups. The development of tools capable of distinguishing closely related languages is crucial for supporting educational and communication needs within Native communities [^2] This study represents a preliminary exploration, and we acknowledge the importance of direct collaboration with Native American communities. Moving forward, we plan to engage with community members and linguistic experts to ensure our work aligns with their perspectives, priorities, and cultural considerations.. We also advocate the decentralization of NLP research efforts, emphasizing the need for targeted investment in endangered languages. Such initiatives are essential to ensure that advances in language technology promote linguistic equity, thereby preserving cultural diversity and heritage in the digital age.
@@
While the study successfully demonstrates the Random Forest classifier’s efficacy in distinguishing Navajo from languages commonly misidentified by Google Translate and identifying related Athabaskan languages, it does have limitations that impact its broader applicability. Firstly, the language variety included in the study is limited; the classifier was tested primarily against a small set of languages suggested by Google’s LangID and a few Athabaskan languages. This narrow scope might not capture the classifier’s effectiveness across a broader range of Native American languages, potentially limiting its utility for other endangered language families. Secondly, the experimental design assumes a binary distinction between Navajo and other languages without considering intra-group variations and dialectical differences within the Athabaskan language family, which could affect accuracy in real-world applications. Lastly, reliance on vectorized features of 5,000 dimensions may overlook some finer linguistic nuances, which are crucial for distinguishing between closely related languages. Addressing these limitations in future work will be essential for developing more robust and applicable language identification systems.
@@
Ethical considerations are paramount in the development of language technology, especially for Native American languages, which are deeply intertwined with cultural identity and heritage. This study emphasizes the importance of respectful engagement with these communities, recognizing the cultural, spiritual, and historical significance of their languages. Technology development involving Native American languages should proceed with close collaboration with native speakers and community leaders to ensure that these tools support and reinforce language preservation rather than contributing to cultural homogenization or appropriation. Additionally, data privacy and consent are critical, as much of the linguistic data involves sensitive cultural content. Ensuring that communities retain control over how their linguistic resources are used is essential for maintaining trust and upholding ethical standards in research. Moreover, to ensure transparency and foster research, we have made our code and datasets used publicly available at https://github.com/ivoryayang/Isitnavajo.
Acknowledgment
This work was partially funded by a Google Research Award, and in part by a Dartmouth Alumni Research Award. We extend our gratitude to the Dartmouth graduate alumni for their generous support and commitment to fostering academic inquiry.